Recent Log Entries

Operator SIGN IN

Log Number: 4004379Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Schaumloffel
DATE:             06/18/2022
SHIFT:               Swing

The operator has read the following:

Monitoring Beam Profile/Active Collimator

Log Number: 4004367Logbook(s): Body: 

Channels: AC:inner:position:x, AC:inner:position:y, bpu_mean_x, bpu_mean_y

Last three hours. Feel free to leave a note if want to see more. Previous conditions "blow up" BPU plots.

1L22 Cryomodule FSD Trip but No Cavity Trip

Log Number: 4004360Logbook(s): Entry Makers: bjacobBody: 

Cavity 7 did not trip (at least to our regular human eyes). GMES did spike before each of these "ghost" FSD trips. Maybe the cavity is just recovering super quick...

It seems to only happen at >50 uA to Hall C.

Autolog save

Log Number: 4004359Logbook(s): Tags: Body: saver ran from opsl05

Saved: moser 06-18-22 13:24:09 {A:2 B:5 C:5 E:1047 A=BS B=CW C=CW D=CW 0R08=56.67 No dumps in.

iCalibrate: Hall C

Log Number: 4004357Logbook(s): Body: Calibration ParametersHall:CTarget: Loop 2 10cmPass: 5Note: 2x2 Raster, HMS: 22*, SHMS -23* Calibrated On:2022-Jun-17 20:07Calibrated By: mmerzFile:C-2022-06-17_2007_Loop_2_10cm_Pass-5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

8E02 Viewer -- C and D Beams

Log Number: 4004347Logbook(s): Entry Makers: moserBody: 

While normally 100 on C-laser atten in VL is too much for most viewers, needed 130 on atten to get Figure 1, 8E02 viewer. Figure 2: D-beam.

Linked log, C-laser harp swipe at 8E01.

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 4004345Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: lrichard
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T

Begin check beam profiler

Log Number: 4004334Logbook(s): Body: 

Channels: bpu_mean_x, bpu_mean_y, bpu:x:total_rate

not much from this plot alone, except some spikes on the x and y. pending more studies

Is 2L03-5 truly stepped off?

Log Number: 4004322Logbook(s): Entry Makers: lrichardBody: 

I'm asking if we can trust that 2L03-5 is truly stepped off because I see the SL lock cavites do some strange tuning at times. If the tuner is broken, does tuning it 1,106.000 steps take it off resonance? Is it the fact that the cavity does not read any gradient that gives us the warm and fuzzy that it is stepped off?

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 4004319Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: lrichard
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T

Pages